Problem Solving Using Systems of Equations
Julie and Eric row their boat(at a constant speed) 55 miles downstream for 5 hours, helped by the current. Row at the same rate, the trip back against the current takes 11 hours. Find the rate of the current.

Answer by Earlsdon(6294)

You can use the distance formula (d = rt) to solve this problem.
Downstream trip: d = 55 miles, t = 5 hours, and r = R + C
R + C is their rowing speed (R) plus the rate of the current (C).
Upstream trip: d = 55 miles, t = 11 hours, and r = R - C
R - C is their rowing speed (R) minus the the rate of the current (C).
Downstream: 55 mi. = (R + C)5 hrs. Divide both sides by 5: R + C = 11
Upstream: 55 mi. = (R - C)11 hrs. Divide both sides by 11: R - C = 5
R + C = 11 Subtract C from both sides.
R = 11 - C Substitute this into: R - C = 5
(11 - C) - C = 5 Simplify and solve for C
11 - 2C = 5 Add 2C to both sides.
11 = 2C + 5 Subtract 5 from both sides.
6 = 2C Divide both sides by 2
C = 3 mph is the rate of the current.